OpenStudy (bry856):

The roots of an equation are x = -1 ± i. The equation is x^2 + _____ + 2 = 0.

OpenStudy (sdfgsdfgs):

"The roots of an equation are x = -1 ± i. " so the eqn is (x-(-1+i))*(x-(-1-i)) = 0 expend the above n simply...what will u get?

jimthompson5910 (jim_thompson5910):

Let p and q be the roots of \(\Large x^2 + bx + c = 0\) Rule: \(\Large p+q = -b\) so \(\Large b = -(p+q)\)

OpenStudy (campbell_st):

well if you think about the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] you can see from the equation a = 1 and c = 2 so \[x = \frac{-b \pm \sqrt{b^2 - 4 \times 1 \times 2}}{2 \times 1}\] then comparing the parts \[-1 = \frac{-b}{2}\] solve for b... and then check with b by using the rest of the quadratic formula \[b^2 - 8 = \frac{\sqrt{4i^2}}{2}\]